OpenStudy (anonymous):

The booster club sold hamburgers and hot dogs at a high school basketball game. They sold 300 hamburgers and 400 hot dogs, and raised a total of $1500. The combined cost of one hamburger and one hotdog was $4. The equations and graph below can be used to determine how much each type of food cost, where m represents the cost of one hamburger and h represents the cost of one hot dog. Total food sales: 300m + 400h = 1500 Total cost for one hamburger and one hot dog: m + h = 4 What was the price, in dollars, of a hotdog?

OpenStudy (jdoe0001):

m + h = 4 <--- can you solve that for say "h" ?

OpenStudy (anonymous):

What do you mean?

OpenStudy (jdoe0001):

isolate "h" on the left-hand side

OpenStudy (anonymous):

Oh! 4 - m = h ?

OpenStudy (jdoe0001):

yeap so , now we know that 4-m = h actually..... simpler if we had isolated "m" but anyhow 4-m = h or one could also say 4-h = m if we isolate "m" so let us use that in the other equation \(\bf m + h = 4\implies 4-h={\color{red}{ m}} \\ \quad \\ 300{\color{red}{ m}} + 400h=1500\implies 300({\color{red}{ 4-h}}) + 400h=1500\) then solve for "h", or isolate "h" to know how much a hamburger costs

OpenStudy (anonymous):

So it's 3?

OpenStudy (anonymous):

Oh my gosh thank you so much!

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

must be really good hotdogs for $3 =)

OpenStudy (anonymous):

haha yes. thanks again by the way!

OpenStudy (jdoe0001):

yw